Find an equation of a line that passes through the points (-1,2) and (5,0)
Thank you

Answer by rfer(16322)

slope
m=(0-2)/(5--1)
m=-2/6
m=-1/3
----------------
y=mx+b
2=-1/3*-1+b
2=1/3+b
b=5/3
------------
y=-1/3x+5/3